OpenStudy (anonymous):

If a sphere has a surface area of 324pi cm^2, find the volume of the sphere

OpenStudy (mertsj):

Use this: A=4 pi r^2 to find r Then use this: V = 4/3 pi r^3 to find the volume

OpenStudy (anonymous):

Can you plug it in? I've literally been trying that formula all night and its not turning out

OpenStudy (anonymous):

324pi=4*3.14 r^2

OpenStudy (mertsj):

Leave pi as pi. Don't substitute 3.14. Divide both sides by 4

OpenStudy (anonymous):

So 324 divide by 4?

OpenStudy (mertsj):

\[r=\frac{9}{\sqrt{\pi}}\]

OpenStudy (mertsj):

Yes/ If you divide both sides by 4 you will have to divide 324 by 4 since it is the left side.

OpenStudy (anonymous):

I got 81

OpenStudy (anonymous):

Does that seem right

OpenStudy (mertsj):